# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

### v2

@@ 8 Conclusion
We conduct a unifying empirical study of parameter-efficient fine-tuning (PETL), an emerging topic in the large model era. We have several new insights and implications, including PETL methods’ complementary expertise, suitable application regimes, and robustness to domain shifts. We expect our study to open new research directions and serve as a valuable user guide in practice.
9 Reproducibility Statement
We have made extensive efforts to ensure the reproducibility of our results. All 14 PETL algorithms and two baseline models, along with data preprocessing routines and data loaders for all datasets—including 19 low-shot datasets, 3 many-shot (full) datasets, and 5 robustness-related datasets—are implemented within a systematic and extensible framework. This framework is designed to facilitate the easy addition of new PETL methods and datasets, modification of backbones, and incorporation of additional scenarios, serving as a convenient tool for future research. Detailed explanations of our implementations, raw results for all experiments, and commands to reproduce the results are thoroughly documented in the README file. We provide the anonymous source code in the supplementary material.
10 Ethics Statement
Our study provides a unifying study of PETL in visual recognition. We expect it to serve as a valuable practical user guide to benefit society. Specifically, fine-tuning large models needs significant computation. A unifying study of PETL will ease end-users to apply more parameter-efficient and computation-efficient ways for fine-tuning. To our knowledge, our paper does not introduce any additional negative societal impacts compared to existing papers on PETL.
@@ A.1 Experiment Details
We employ AdamW optimizer Loshchilov & Hutter (2018) with a batch size of 64 and utilize the cosine decay learning rate scheduler. We train all methods with 100 epochs. The learning rate is tuned from [1e-3, 1e-2] and weight decay from [1e-4, 1e-3]. The method-specific hyperparameter searching grip is shown in Table 3, along with the tunable parameter ranges (in millions). Since most method-specific hyperparameters affect the number of tunable parameters in the PETL methods, we set a cap on the tunable parameters for each PETL method to be less than or equal to 1.5% of the total parameters in ViT-B/16, which is approximately equal to the number of parameters in the Query, Key, and Value matrices of a single MSA block. Consistent with the original VTAB-1k paper (Zhai et al., 2019), most PETL studies Jie & Deng (2023; 2022); Luo et al. (2023); Jia et al. (2022); Zhang et al. (2022b); Luo et al. (2023); Lian et al. (2022) don’t apply data augmentation as it’s challenging to identify a set of augmentations that uniformly benefits all 19 datasets [^4] To demonstrate it, we apply simple data augmentations (RandomResizedCrop and RandomHorizontalFlip) on three datasets in each group, as shown in Table 4. . To ensure that our results are directly comparable and that any performance differences are attributable to the methods themselves rather than data augmentation, we don’t apply data augmentation.
@@
In section 3 and section 5, the fine-tuning train set and test set are from the same data distribution. In section 7, we fine-tune the CLIP model using ImageNet-1K training data (100 shots) and subsequently evaluate the fine-tuned model not only on the test set of ImageNet-1K but also on four additional datasets with distribution shifts: ImageNet-V2, ImageNet-R, ImageNet-S, and ImageNet-A, as shown in Figure 8. Following Wortsman et al. (2022), we set a small learning rate as 3e-5 and weight decay as 5e-3. We use a strong data augmentation following Zhang et al. (2022b).
(a) ImageNet Deng et al. (2009)
(b) ImageNetV2 Recht et al. (2019)
(c) ImageNet-R Hendrycks et al. (2021a)
(d) ImageNet Sketch Wang et al. (2019)
(e) ImageNet-A Hendrycks et al. (2021b)
Figure 8: Samples of the class lemon, from the fine-tuned dataset ImageNet and distribution shifts datasets (ImageNet-V2, ImageNet-R, ImageNet-S, and ImageNet-A). The CLIP model is fine-tuned with PETL on ImageNet and evaluated on distribution shifts datasets to measure the robustness of fine-tuned models. The figures are modified based on Wortsman et al. (2022).
@@ A.2 Dataset Details
Table 4: We apply simple data augmentations (DA) (RandomResizedCrop and RandomHorizontalFlip) on three datasets in each group. Data augmentation generally does not benefit VTAB-1K and thus, most recent PETL papers (Jie & Deng, 2023; 2022; Luo et al., 2023; Jia et al., 2022; Zhang et al., 2022b; Luo et al., 2023; Lian et al., 2022) skip it. To ensure that our results are directly comparable to existing papers, we don’t apply data augmentation.
@@ B.1 Vision Transformer
Overview of ViT. Inspired by the recent success of Transformer-based models Vaswani et al. (2017) in NLP Wolf et al. (2020), Vision Transformer (ViT) Dosovitskiy et al. (2020) has become widely used in computer vision. To handle 2D images, ViT divides an image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ non-overlapping patches $\{\bm{I}^{(n)}\in\mathbb{R}^{P^{2}\times C}\}_{n=1}^{N}$, where $(H,W)$ is the resolution of the input image, $C$ is the number of channels, $N=HW/P^{2}$ and $(P,P)$ is the resolution of each patch. Each patch $\bm{I}^{(n)}$ is flattened and embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with a trainable linear projection. Incorporating the BERT design approach Kenton & Toutanova (2019), a “Class” token ${\bm{x}}_{0}^{(\text{Class})}$ is prepended to the sequence of embedded patches, whose output state at the last Transformer layer is utilized as the image representation. Finally, position embeddings $\mathbf{E}_{\text{pos }}\in\mathbb{R}^{{D\times(1+N)}}$ are added to preserve positional information and form the input $\mathbf{Z}_{0}\in\mathbb{R}^{D\times(1+N)}$ to the ViT, which can be formulated by:
@@
Details of each Transformer layer. As shown in the right part of Figure 9, each Transformer layer consists of a Multi-head Self-Attention (MSA) block, a Multi-Layer Perceptron (MLP) block, and two Layer Normalization (LN) layers Ba et al. (2016). Formally, a Transformer layer $L_{m}$ can be defined as
@@
Since PETL methods often entail incorporating additional components to modify the intermediate features within or between Transformer layers, we adopt the notation $\{h_{1},\ldots,h_{10}\}$ to denote the intermediate features in the unravelled view of a Transformer layer (as depicted in Figure 9) to facilitate a clearer illustration of the PETL methods discussed in the subsequent section.
@@ B.2 Evaluated Methods
In this section, we dive into the details of 12 State-Of-The-Art PETL approaches, categorized into three groups: Prompt-based, Adapter-based and Selective Parameter Tuning. We will describe the distinctions and tradeoffs between them. A consolidated overview of these approaches is summarized in Table 6.
Figure 9: An overview of a Transformer block in ViT. We adopt the notation $\{h_{1},\ldots,h_{10}\}$ to denote the intermediate features within a Transformer block to facilitate a clearer illustration of the PETL methods discussed in subsection B.2.
Table 6: PETL Methods Summary: Prompt-based and adapter-based methods incorporate additional parameters to modify features while keeping the original backbone intact. However, these added parameters introduce additional inference overhead. In contrast, selective tuning methods modify the backbone by updating selective parameters, thereby incurring no additional inference overhead.
@@ B.2.1 Prompt-based Methods
Prompt-based learning emerged in NLP as an effective approach to adapt pre-trained models for downstream tasks Liu et al. (2023); Lialin et al. (2023). The core concept involves augmenting the model input with task-specific hints (prompts), which aid the pre-trained model in addressing novel tasks with its existing knowledge. Hard prompts are human-interpretable natural language hints, encompassing task instructions, in-context examples, or supporting information. Alternatively, soft prompts are continuous vector hints that are incorporated into the input embeddings of the input layers or hidden states of other layers. Soft prompts are updated during the fine-tuning process using gradient-based methods, guided by the downstream task-specific loss functions, while the pre-trained model itself remains fixed. The splendent success of prompts in NLP has sparked a growing interest in adopting it in computer vision Yu et al. (2023b); Tu et al. (2023) and multi-modal domains Gu et al. (2023).
In this paper, we investigate a prominent and strong prompt-based method called Visual Prompt Tuning (VPT) Jia et al. (2022), which represents one of the early endeavours in introducing prompts to computer vision. Specifically, VPT-Shallow adds $l$ prompts $\bm{P}_{0}\in\mathbb{R}^{l\times D}$ to the input of the first Transformer layer $\bm{Z}_{0}$ and the output $\tilde{\bm{P}_{0}}$ of $\bm{P}_{0}$ serves as the input for the next layer as depicted in Equation 11. VPT-Shallow can be perceived as the addition of learnable pixels to the original images. On the other hand, VPT-Deep inserts $l$ prompts $\{\bm{P}_{m}\in\mathbb{R}^{l\times D}\}_{m=0}^{M}$ to the input of every Transformer layer $\bm{Z}_{m}$ but their outputs are discarded at the end of the layer as illustrated in Equation 12.
@@ B.2.2 Adapter-based Methods
Adapter-based methods typically introduce additional trainable parameters into a frozen pre-trained model to facilitate learning of downstream tasks Lialin et al. (2023). Initially developed for multi-domain adaptation Rebuffi et al. (2017; 2018) and continual learning Rosenfeld & Tsotsos (2018); Mai et al. (2022), the idea of Adapters is subsequently embraced by Houlsby et alHoulsby et al. (2019) in the NLP domain to adapt Transformer-based networks for downstream tasks, and it also has garnered increasing interest in the computer vision field Yu et al. (2023b). In this comparative analysis, we concentrate on five popular Adapter-based methods, encompassing the original Adapter, along with variants focusing on adjusting the positions of Adapters Chen et al. (2022b); Pfeiffer et al. (2021), introducing visual inductive biases Jie & Deng (2022), as well as employing re-parameterization to reduce the number of trainable parameters and inference latency Luo et al. (2023).
@@
Houlsby et al. (2019) inserts two lightweight bottleneck-structured modules into each Transformer layer: one after the MSA block and the other after the MLP block. As depicted in 10(a), the Adapter is composed of a down-projection layer with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a nonlinear activation function $\sigma$, an up-projection layer with $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, a scaling factor $s$ and a skip-connection. To limit the number of trainable parameters, the bottleneck dimension is much smaller than the feature dimension $r\ll D$. Formally, Houl. Adapter can be defined as:
@@
Pfeiffer et al. (2021) is a more efficient variant that introduces the Adapter solely after the MLP block, a strategy that has demonstrated effectiveness in recent studies Hu et al. (2021). Pfeif. Adapter can be defined formally as $h_{9}=\operatorname{Adapter}(h_{9})$ where $\operatorname{Adapter}$ follows Equation 14.
@@
Chen et al. (2022b) proposed to insert the Adapter in parallel with the MLP block, which differs from the sequential design of Houl. and Pfeif. Adapter. The rationale behind this parallel design lies in the belief that the domain-specific features generated by the Adapter can complement the domain-agnostic features derived from the original MLP block, leading to an improved feature ensemble Szegedy et al. (2015). Formally, AdaptFormer can be defined as $\bm{h_{9}}=\bm{h_{9}}+\operatorname{Adapter}\bm{(h7)}$ where $\operatorname{Adapter}$ follows Equation 14.
@@
(Convolutional By-Passes) Jie & Deng (2022) addresses the concern that many existing Adapters lack visual inductive bias, potentially limiting their performance for downstream vision tasks with limited data. To this end, the authors introduce a convolutional bottleneck module, running in parallel with the MSA or(and) MLP block. This module encompasses a $1\times 1$ convolution reducing the channel with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a $3\times 3$ convolution with the same input and output channel, a $1\times 1$ convolution expanding the channel $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, two nonlinear functions $\sigma$ and a scaling factor $s$, as shown in 10(b). The authors argue that Convpass is more efficient at capturing visual information in low-data scenarios due to its hard-coded locality of convolutional layers. The formal definition of Convpass is shown in Equation 15.
@@
Luo et al. (2023) found that the removal of the nonlinear function in the Adapter does not result in performance degradation for vision tasks. In light of this finding, the authors propose a linear Adapter with group-wise transformation Luo et al. (2022) and sequentially added two of these linear Adapters to both MSA and MLP blocks. Owing to the sequential placement of the RepAdapter and its inherent linearity, the additional parameters can be re-parameterized to the original MSA or MLP block after training, thereby incurring zero additional costs during inference. RepAdapter is illustrated in 10(c) and formally defined in Equation 16.
@@
where $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, $\tilde{h}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{r}{G}\times(N+1)}$ is the features splitted from $\tilde{h}\in\mathbb{R}^{r\times(N+1)}$ and $G$ is the number of groups in group-wise transformation Luo et al. (2022). $\bm{W}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{D}{G}\times\frac{r}{G}}$ is the projection weight matrix.

### v4

@@ 8 Conclusion
Instead of chasing the leaderboard, we conduct a unifying empirical study of PEFT, an emerging topic in the large model era. We provide an extendable framework for reproducible evaluations of PEFT methods in computer vision. We also have several new insights and implications, including PEFT methods’ complementary expertise, suitable application regimes, and robustness to domain shifts. We expect our study to open new research directions and serve as a valuable user guide in practice.
Acknowledgment
This research is supported by grants from the National Science Foundation (ICICLE: OAC-2112606). We are grateful for the generous support from the Google gift fund and the Ohio Supercomputer Center.
@@ A.1 Experiment Details
We employ AdamW optimizer [61] with a batch size of 64 and utilize the cosine decay learning rate scheduler. We train all methods with 100 epochs. The learning rate is tuned from [1e-3, 1e-2] and weight decay from [1e-4, 1e-3]. The method-specific hyperparameter searching grip is shown in Table 3, along with the tunable parameter ranges (in millions). Since most method-specific hyperparameters affect the number of tunable parameters in the PEFT methods, we set a cap on the tunable parameters for each PEFT method to be less than or equal to 1.5% of the total parameters in ViT-B/16, which is approximately equal to the number of parameters in the Query, Key, and Value matrices of a single MSA block. Consistent with the original VTAB-1k paper [104], most PEFT studies [44, 43, 63, 42, 107, 63, 55] don’t apply data augmentation as it’s challenging to identify a set of augmentations that uniformly benefits all 19 datasets [^4] To demonstrate it, we apply simple data augmentations (RandomResizedCrop, RandomVerticalFlip and RandomHorizontalFlip) on three datasets in each group, as shown in Table 4. . To ensure that our results are directly comparable and that any performance differences are attributable to the methods themselves rather than data augmentation, we don’t apply data augmentation.
@@
In section 3 and section 5, the fine-tuning train set and test set are from the same data distribution. In section 7, we fine-tune the CLIP model using ImageNet-1K training data (100 shots) and subsequently evaluate the fine-tuned model not only on the test set of ImageNet-1K but also on four additional datasets with distribution shifts: ImageNet-V2, ImageNet-R, ImageNet-S, and ImageNet-A, as shown in Figure 7. Following [96], we set a small learning rate as 3e-5 and weight decay as 5e-3. We use a strong data augmentation following [107].
(a) ImageNet [15]
(b) ImageNetV2 [78]
(c) ImageNet-R [34]
(d) ImageNet Sketch [93]
(e) ImageNet-A [35]
Figure 7: Samples of the class lemon, from the fine-tuned dataset ImageNet and distribution shifts datasets (ImageNet-V2, ImageNet-R, ImageNet-S, and ImageNet-A). The CLIP model is fine-tuned with PEFT on ImageNet and evaluated on distribution shifts datasets to measure the robustness of fine-tuned models. The figures are modified based on [96].
@@
What do we not investigate?
There are many aspects that one can ask about PEFT. Our study focuses more on their learning and prediction behaviors, not the computation-specific properties like memory usage and FLOPS.
@@ A.2 Dataset Details
Table 4: We apply simple data augmentations (DA) (RandomResizedCrop, RandomVerticalFlip and RandomHorizontalFlip) on three datasets in each group. Data augmentation does not benefit most of VTAB-1K datasets and thus, most recent PEFT papers [44, 43, 63, 42, 107, 63, 55] skip it. Figure 8 shows examples of how some data augmentation transforms are harmful for a specific task. Therefore, to ensure that our results are directly comparable to existing papers, we don’t apply data augmentation.
(a) Original
(b) Vertical Flipped
(c) Cropped
(d) Resized
Figure 8: Example of augmented images from KITTI: (a) Original, (b) RandomVerticalFlip, (c) RandomResizedCrop, (d) Resize only. The KITTI task needs to predict the depth to the nearest vehicle (car, van, or truck) in the image. RandomResizedCrop may crop out the nearest vehicle. RandomVerticalFlip may make the task more difficult.
@@ B.1 Vision Transformer
Overview of ViT. Inspired by the recent success of Transformer-based models [90] in NLP [95], Vision Transformer (ViT) [19] has become widely used in computer vision. To handle 2D images, ViT divides an image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ non-overlapping patches $\{\bm{I}^{(n)}\in\mathbb{R}^{P^{2}\times C}\}_{n=1}^{N}$, where $(H,W)$ is the resolution of the input image, $C$ is the number of channels, $N=HW/P^{2}$ and $(P,P)$ is the resolution of each patch. Each patch $\bm{I}^{(n)}$ is flattened and embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with a trainable linear projection. Incorporating the BERT design approach [46], a “Class” token ${\bm{x}}_{0}^{(\text{Class})}$ is prepended to the sequence of embedded patches, whose output state at the last Transformer layer is utilized as the image representation. Finally, position embeddings $\mathbf{E}_{\text{pos }}\in\mathbb{R}^{{D\times(1+N)}}$ are added to preserve positional information and form the input $\mathbf{Z}_{0}\in\mathbb{R}^{D\times(1+N)}$ to the ViT, which can be formulated by:
@@
Details of each Transformer layer. As shown in the right part of Figure 9, each Transformer layer consists of a Multi-head Self-Attention (MSA) block, a Multi-Layer Perceptron (MLP) block, and two Layer Normalization (LN) layers [4]. Formally, a Transformer layer $L_{m}$ can be defined as
@@
Since PEFT methods often entail incorporating additional components to modify the intermediate features within or between Transformer layers, we adopt the notation $\{h_{1},\ldots,h_{10}\}$ to denote the intermediate features in the unravelled view of a Transformer layer (as depicted in Figure 9) to facilitate a clearer illustration of the PEFT methods discussed in the subsequent section.
@@ B.2 Evaluated Methods
In this section, we dive into the details of 12 State-Of-The-Art PEFT approaches, categorized into three groups: Prompt-based, Adapter-based and Selective Parameter Tuning. We will describe the distinctions and tradeoffs between them. A consolidated overview of these approaches is summarized in Table 6.
Figure 9: An overview of a Transformer block in ViT. We adopt the notation $\{h_{1},\ldots,h_{10}\}$ to denote the intermediate features within a Transformer block to facilitate a clearer illustration of the PEFT methods discussed in subsection B.2.
Table 6: PEFT Methods Summary: Prompt-based and adapter-based methods incorporate additional parameters to modify features while keeping the original backbone intact. However, these added parameters introduce additional inference overhead. In contrast, selective tuning methods modify the backbone by updating selective parameters, thereby incurring no additional inference overhead.
@@ B.2.1 Prompt-based Methods
Prompt-based learning emerged in NLP as an effective approach to adapt pre-trained models for downstream tasks [57, 54]. The core concept involves augmenting the model input with task-specific hints (prompts), which aid the pre-trained model in addressing novel tasks with its existing knowledge. Hard prompts are human-interpretable natural language hints, encompassing task instructions, in-context examples, or supporting information. Alternatively, soft prompts are continuous vector hints that are incorporated into the input embeddings of the input layers or hidden states of other layers. Soft prompts are updated during the fine-tuning process using gradient-based methods, guided by the downstream task-specific loss functions, while the pre-trained model itself remains fixed. The splendent success of prompts in NLP has sparked a growing interest in adopting it in computer vision [102, 88] and multi-modal domains [27].
In this paper, we investigate a prominent and strong prompt-based method called Visual Prompt Tuning (VPT) [42], which represents one of the early endeavours in introducing prompts to computer vision. Specifically, VPT-Shallow adds $l$ prompts $\bm{P}_{0}\in\mathbb{R}^{l\times D}$ to the input of the first Transformer layer $\bm{Z}_{0}$ and the output $\tilde{\bm{P}_{0}}$ of $\bm{P}_{0}$ serves as the input for the next layer as depicted in Equation 11. VPT-Shallow can be perceived as the addition of learnable pixels to the original images. On the other hand, VPT-Deep inserts $l$ prompts $\{\bm{P}_{m}\in\mathbb{R}^{l\times D}\}_{m=0}^{M}$ to the input of every Transformer layer $\bm{Z}_{m}$ but their outputs are discarded at the end of the layer as illustrated in Equation 12.
@@ B.2.2 Adapter-based Methods
Adapter-based methods typically introduce additional trainable parameters into a frozen pre-trained model to facilitate learning of downstream tasks [54]. Initially developed for multi-domain adaptation [76, 77] and continual learning [80, 66], the idea of Adapters is subsequently embraced by Houlsby et al. [36] in the NLP domain to adapt Transformer-based networks for downstream tasks, and it also has garnered increasing interest in the computer vision field [102]. In this comparative analysis, we concentrate on five popular Adapter-based methods, encompassing the original Adapter, along with variants focusing on adjusting the positions of Adapters [11, 74], introducing visual inductive biases [43], as well as employing re-parameterization to reduce the number of trainable parameters and inference latency [63].
@@
[36] inserts two lightweight bottleneck-structured modules into each Transformer layer: one after the MSA block and the other after the MLP block. As depicted in 10(a), the Adapter is composed of a down-projection layer with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a nonlinear activation function $\sigma$, an up-projection layer with $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, a scaling factor $s$ and a skip-connection. To limit the number of trainable parameters, the bottleneck dimension is much smaller than the feature dimension $r\ll D$. Formally, Houl. Adapter can be defined as:
@@
[74] is a more efficient variant that introduces the Adapter solely after the MLP block, a strategy that has demonstrated effectiveness in recent studies [37]. Pfeif. Adapter can be defined formally as $h_{9}=\operatorname{Adapter}(h_{9})$ where $\operatorname{Adapter}$ follows Equation 14.
@@
[11] proposed to insert the Adapter in parallel with the MLP block, which differs from the sequential design of Houl. and Pfeif. Adapter. The rationale behind this parallel design lies in the belief that the domain-specific features generated by the Adapter can complement the domain-agnostic features derived from the original MLP block, leading to an improved feature ensemble [85]. Formally, AdaptFormer can be defined as $\bm{h_{9}}=\bm{h_{9}}+\operatorname{Adapter}\bm{(h7)}$ where $\operatorname{Adapter}$ follows Equation 14.
@@
(Convolutional By-Passes) [43] addresses the concern that many existing Adapters lack visual inductive bias, potentially limiting their performance for downstream vision tasks with limited data. To this end, the authors introduce a convolutional bottleneck module, running in parallel with the MSA or(and) MLP block. This module encompasses a $1\times 1$ convolution reducing the channel with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a $3\times 3$ convolution with the same input and output channel, a $1\times 1$ convolution expanding the channel $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, two nonlinear functions $\sigma$ and a scaling factor $s$, as shown in 10(b). The authors argue that Convpass is more efficient at capturing visual information in low-data scenarios due to its hard-coded locality of convolutional layers. The formal definition of Convpass is shown in Equation 15.
@@
[63] found that the removal of the nonlinear function in the Adapter does not result in performance degradation for vision tasks. In light of this finding, the authors propose a linear Adapter with group-wise transformation [62] and sequentially added two of these linear Adapters to both MSA and MLP blocks. Owing to the sequential placement of the RepAdapter and its inherent linearity, the additional parameters can be re-parameterized to the original MSA or MLP block after training, thereby incurring zero additional costs during inference. RepAdapter is illustrated in 10(c) and formally defined in Equation 16.
@@
where $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, $\tilde{h}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{r}{G}\times(N+1)}$ is the features splitted from $\tilde{h}\in\mathbb{R}^{r\times(N+1)}$ and $G$ is the number of groups in group-wise transformation [62]. $\bm{W}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{D}{G}\times\frac{r}{G}}$ is the projection weight matrix.